Given demand equation P=50-2Qd and supply equation P=10+2Qs calculate. 

(a) consumer surplus

(b) producer surplus

(c) total surplus

assuming pure competition


1
Expert's answer
2021-06-02T11:58:57-0400

"P=50-2Qd\\P=10+2Qs\\50-2Q=10+2Q\\50-10=2Q+2Q\\40=4Q\\Q=10"


"P=10+2Q\\=10+92\times10)\\P=30"


consumer surplus

"\frac{1}{2}\times base \times height"

We calculate Pd when Qd=0

"P=50-2Qd"


"\frac{2Qd}{2}=\frac{50}{2}-\frac{P}{2}"


"Qd=25-\frac{1}{2}P"


"0=25-\frac{1}{2}P"


"P=50"


consumer surplus

"=\frac{1}{2}\times10\times(50-30)"

"=\frac{1}{2}\times10\times20\\=100"


producer surplus

"P=10+2Qs"

"\frac{2Qs}{2}=\frac{P}{2}-\frac{10}{2}"

"Qs=\frac{1}{2}P-5"

"0=\frac{1}{2}P-5"

"P=10"


"\frac{1}{2}\times10\times(30-10)\\=100"


Total surplus

consumer surplus + producer surplus

"=100+100\\=200"